# On the conjugates of Christoffel words

### v4

@@ 8. Borders of conjugates of Christoffel words
8.1. Borders
In this subsection, we determine the longest border of every conjugate of a Christoffel word, thereby reproving a result of Lapointe [La] (but with a totally different method). Indeed, the length of the longest border and the smallest nontrivial period of a word are related: their sum is the length of the word.
@@ 8.1. Borders
The key point for the proof of Theorem 8.1 is the determination of all the occurrences of $V_{m-1}$ in $V_{m}$. Exactly $b_{m}$ of them can be read on the factorization (8.1), but there may be additional ones. By primitivity of $V_{m-1}$, the word $V_{m-1}V_{m-1}$ contains exactly two occurrences of $V_{m-1}$. Consequently, if an additional occurrence of $V_{m-1}$ appears, then it must be a factor of $V_{m-1}V_{m-2}V_{m-1}$. A more precise statement is given in Lemma 8.14.
@@ 8.2. Consequences
As noted by one of the referees, this result may also be obtained as follows: if $u$ is a border of the Christoffel word $w$, then $uu$ is a factor of $ww$; since $ww$ is a Sturmian word, it is Sturmian, and thus $uu$ too. Hence all conjugates of $u$ are Sturmian. Among them is the power $\ell^{k}$ of some Lyndon word $\ell$, which is therefore Sturmian; hence $\ell$ is a Christoffel word by a theorem of Berstel and de Luca (see [Re] Corollary 13.4.3).
@@
where $M_{m}=V_{m}(0,\ldots,0)$ is as before the word corresponding to the sequence $a_{1},\ldots,a_{m}$ and to the Ostrowski representation of 0. Note that here $N$ may be in $\mathbb{Z}$; but the word $H_{N}(a_{1},\ldots,a_{m})$ depends only on $N\mod q_{m}$, where $q_{m}$ is the length of this word (recall that $q_{j}=K_{j}(a_{1},\ldots,a_{j})$).
@@
Note that we recover a step function (the number $t$ in the statement) as it appears in Lapointe’s article, see for example [La, Figure 3].
@@
Proof.
Denote the new sequence by $a^{\prime}_{1}=a_{2}+1,a^{\prime}_{2}=a_{3},\ldots,a^{\prime}_{m-1}=a_{m}$; then the associated sequence of $b_{i}$’s is $b^{\prime}_{1}=a^{\prime}_{1}-1=a_{2},b^{\prime}_{2}=a^{\prime}_{2},\ldots,b^{\prime}_{m-1}=a^{\prime}_{m-1}$. Recall that $V_{-1}=V^{\prime}_{-1}=b,V_{0}=V^{\prime}_{0}=a$. Observe that $V_{1}=b$, $V_{2}=b^{a_{2}-d_{2}}ab^{d_{2}}$, while $V^{\prime}_{1}=a^{a_{2}-d_{2}}ba^{d_{2}}$. Thus, $V_{1}=E(V^{\prime}_{0})$ and $V_{2}=E(V^{\prime}_{1})$. An immediate induction based on (7.1) proves the lemma. ∎
@@
Proof.
Let $q_{i}$ be as usual and write $q^{\prime}_{i}$ for the corresponding numbers with respect to the sequence $a_{1},\ldots,a_{m-2},a_{m-1}+1$. By stability, $V_{i}=V^{\prime}_{i}$ for $i=1,\ldots,m-2$. Next $V_{m-1}=V_{m-2}^{b_{m-1}-d_{m-1}}V_{m-3}V_{m-2}^{d_{m-1}}$. Thus,
$V_{m}(d_{1},\ldots,d_{m-1},0)=V_{m-1}V_{m-2}$ $=V_{m-2}^{b_{m-1}-d_{m-1}}V_{m-3}V_{m-2}^{d_{m-1}+1}$
$={V^{\prime}}_{m-2}^{b_{m-1}-d_{m-1}}V^{\prime}_{m-3}{V^{\prime}}_{m-2}^{d_{m-1}+1}$
$=V^{\prime}_{m-1}(d_{1},\ldots,d_{m-2},d_{m-1}+1).$
@@
$={V^{\prime}}_{m-2}^{b_{m-1}-d_{m-1}+1}V^{\prime}_{m-3}{V^{\prime}}_{m-2}^{d_{m-1}}$
$=V^{\prime}_{m-1}(d_{1},\ldots,d_{m-2},d_{m-1}),$
and the proof is complete. ∎
Proof of Corollary 8.3.
@@
2. We study now the number $t$ in Part (a) of the statement. Observe that $N=d_{m}q_{m-1}+\sum_{1\leq i\leq m-1}d_{i}q_{i-1}$, so that $\lfloor\frac{N}{q_{m-1}}\rfloor=d_{m}$ by Lemma 10.2 (i).
@@
We show that the numerator in the latter fraction is always in the interval $q_{m-1},q_{m-1})$, so that the integer part of this fraction is either -1 or 0, and we give the condition when it is 0. We have indeed $q_{m-1}\leq q_{m-2}-2-j$, since $j\leq q_{m-1}-1$ by Lemma 10.2 (i), so that $j+2\leq q_{m-1}+1\leq q_{m-1}+q_{m-2}$. Moreover, $q_{m-2}-2-j<q_{m-1}$, since the sequence $q_{i}$ is increasing. Also, if $d_{m-1}>0$, then $j\geq q_{m-2}$, hence $q_{m-2}-2-j<0$; and if $d_{m-1}=0$, then $j=\sum_{1\leq i\leq m-2}d_{i}q_{i-1}$, and $q_{m-2}-2-j<0$ if and only if $j\geq q_{m-2}-1$, which, by Lemmas 10.1 (i) and 10.2 (i), is equivalent to the fact that the sequence $d_{1},\ldots,d_{m-2},0$ is alternating.
@@
Suppose now that $q_{m-1}\leq N<q_{m}$. Then $d_{m}>0$ by Lemma 10.2 (i). We are therefore in case (i), (ii), (iii) or (iv) of Theorem 8.1. Note that $V_{m-1}=H_{J}(a_{1},\ldots,a_{m-1})$, where $J=\sum_{1\leq i\leq m-1}d_{i}q_{i-1}$, so that $V_{m-1}=H_{N}(a_{1},\ldots,a_{m-1})$, because $N\equiv J\mod q_{m-1}$, the length of $V_{m-1}$. Thus we have to show that $B_{N}=V_{m-1}^{t}$.
@@
5. We assume that $a_{m}=1$. Define $a^{\prime}_{i}=a_{i}$ if $i=1,\ldots,m-2$ and $a^{\prime}_{m-1}=a_{m-1}+1$. We denote by $q^{\prime}_{i}$ and $M^{\prime}_{i}$ the corresponding words and numbers. We have $q^{\prime}_{i}=q_{i}$ for $i=1,\ldots,m-2$ and by (2.2), $q^{\prime}_{m-1}=q_{m-1}+q_{m-2}=q_{m}$.
@@
Suppose that $0\leq N<q_{m-1}-1$. Then $q^{\prime}_{m-2}=q_{m-2}\leq N^{\prime}<q_{m-1}+q_{m-2}-1=q_{m}-1=q^{\prime}_{m-1}-1$. It follows from Case (a) that $B_{N}=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})^{t}$, with $t=\min\{\lfloor\frac{N^{\prime}}{q^{\prime}_{m-2}}\rfloor,1+\lfloor\frac{q^{\prime}_{m-1}-2-N^{\prime}}{q^{\prime}_{m-2}}\rfloor\}$. Note that $H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})=C^{N^{\prime}}(H_{0}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2}))=C^{N+q_{m-2}}(H_{0}(a_{1},\ldots,a_{m-2}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-2}))$ (since the word is of length $q_{m-2}$) $H_{N}(a_{1},\ldots,a_{m-2})$; moreover, $t=\min\{\lfloor\frac{N+q_{m-2}}{q_{m-2}}\rfloor,1+\lfloor\frac{q_{m-1}+q_{m-2}-2-N-q_{m-2}}{q_{m-2}}\rfloor\}$, which settles this case.
Suppose now that $q_{m-1}\leq N<q_{m}-1$. Then $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$, where $N^{\prime\prime}=N^{\prime}-q_{m}=N+q_{m-2}-q_{m}=N-q_{m-1}$, since the words have length $q_{m}$. Now $0\leq N^{\prime\prime}<q_{m-2}-1$. Hence by the first part, $B_{N}=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1}-1)=C^{N^{\prime\prime}}(H_{0}(a_{1},\ldots,a_{m-1}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-1}))$ (since the word has length $q_{m-1}$) $H_{N}(a_{0},\ldots,a_{m-1})$. ∎
8.3. Proof of Theorem 8.1
@@
Since $q_{j}$ is the length of $V_{j}$, this proves the lemma. ∎
@@
Proof.
The expression of $X_{k}W_{k}$ is an immediate consequence of Lemma 8.6. Recall Pirillo’s theorem: if the words $aub,bua$ are conjugate, then $u$ is a central word ([P], [Re, Theorem 15.2.5]). By Lemma 8.6, the words $aX_{k}W_{k}b$ and $bX_{k}W_{k}a$ are conjugate. This proves the lemma. ∎
@@
Proof.
If $V_{k+1}$ is a prefix of $V_{k}V_{k+1}$, then the common prefix $W_{k}$ of $V_{k}V_{k+1}$ and $V_{k+1}V_{k}$ is of length $w_{k}\geq q_{k+1}$; since $w_{k}+x_{k}=q_{k+1}+q_{k}-2$, we obtain $x_{k}\leq q_{k}-2$. By Lemma 8.6 this implies that $d_{k+1}=0$. Similarly, if $V_{k+1}$ is a suffix of $V_{k+1}V_{k}$, then $w_{k}\leq q_{k}-2$ and $b_{k+1}-d_{k+1}=0$. ∎
@@
Suppose that the representation (8.2) is greedy. The word $V_{m-1}(d_{1},\ldots,d_{m-1})$ is a prefix of $V_{m}(d_{1},\ldots,d_{m})$ if and only if $V_{m}(d_{1},\ldots,d_{m})$ is not the Christoffel word $V_{m}(\ldots,0,b_{m-2},0,b_{m})$.
Proof.
@@
The lemma then follows from Lemma 3.2 and Corollary 7.4. ∎
@@
Suppose that the representation (8.2) is greedy. The word $V_{m-1}$ is not a prefix of $V_{m-2}V_{m-1}$ if and only if $d_{m-1}=0$ and the sequence $d_{1},\ldots,d_{m-1}$ is alternating.
Proof.
We apply Corollary 8.9 to the sequence $a_{1},\ldots,a_{m-1},1$ and the words $V_{m-1}(d_{1},\ldots,d_{m-1})$ and $V^{\prime}_{m}=V_{m}(d_{1},\ldots,d_{m-1},1)$, so that $V^{\prime}_{m}=V_{m-2}V_{m-1}$: thus the word $V_{m-1}$ is a prefix of $V_{m-2}V_{m-1}$ if and only if $V^{\prime}_{m}$ is not a Christoffel word; but, by Corollary 7.4, $V^{\prime}_{m}$ is a Christoffel word if and only if the sequence $d_{1},\ldots,d_{m-1},1$ is alternating; this means that $d_{m-1}=0$ and the sequence $d_{1},\ldots,d_{m-1}$ is alternating. ∎
@@
Proof.
The word $Y^{i-1}X$ is a border of $Y^{i}X$. Suppose that $B$ is longer. It begins by $Y^{i-1}X$, hence by $Y$ since $i\geq 2$: $B=YB^{\prime}$. Moreover, $Y^{i}X=UB$, where the length of $U$ satisfies $0<|U|<|Y|$. Thus $Y^{i}X=UYB^{\prime}$ and we see that $Y$ is an internal factor of $YY$, a contradiction. ∎
@@
(ii) Follows from (i) and Lemma 8.11. ∎
@@
In the following lemmas, we consider the legal representation (8.2) and put $V_{m}=V_{m}(d_{1},\ldots,d_{m})$, $V_{m-1}=V_{m-1}(d_{1},\ldots,d_{m-1})$, and $V_{m-2}=V_{m-2}(d_{1},\ldots,d_{m-2})$.
@@
The proof for $V_{m-1}V_{m}$ is similar and we omit it. ∎
@@
1. Assume first that $1\leq d_{m}\leq b_{m}-1$. Thus by (8.1), the word $V_{m-1}$ is a prefix and a suffix of $V_{m}$. We show that either $X$ is an integer power of $V_{m-1}$, or $Y$ is an integer power of $V_{m-1}$. Indeed, if $|X|<|V_{m-1}|$, then $X$ is a nontrivial proper suffix of $V_{m-1}$, $V_{m-1}=UX$, where $U$ is nonempty, and $V_{m}$ begins with $XV_{m-1}$; but $V_{m}$ also begins with $V_{m-1}$, $V_{m}=V_{m-1}W$, hence $UV_{m-1}W=UV_{m}=UXV_{m-1}Y=V_{m-1}V_{m-1}Y$; since $U$ is nonempty and shorter than $V_{m-1}$, we see that $V_{m-1}$ is a proper factor of $V_{m-1}^{2}$, and we have a contradiction with the primitivity of $V_{m-1}$.
@@
Suppose now that there is an occurrence of $V_{m}$ ending at $V_{m-1}$; this is equivalent to the fact that $V_{m}$ is a suffix of $V_{m}V_{m-1}$. Then, similarly, we must have $d_{m}=b_{m}$. ∎
@@
Suppose that $B$ is longer; then $B$ extends $Y^{i}$ to the left, hence $B$ ends by $Y$, since $i\geq 1$; moreover, we may extend the prefix $Y^{i}$ of $W$ to the longer prefix $B$, and since $B$ ends by $Y$, by 1. this $Y$ is the factor $Y$ of $W$ starting at $X$, and we are in case L; thus, since $B$ ends by $Y$, by 1. and by the hypothesis on the locations of the factors $Y$ in $W$, $B=Y^{i+1}$. But $B$ is also a right factor of $W$, hence we must have $j>i$, otherwise there is a factor $Y$ ending at $X$, which is excluded by 2. ∎
@@
Then, since $B$ is a suffix of $XY^{j}$, we have $XY^{j}=W{\boldsymbol{Y}}UZ$; then we see that that either $\boldsymbol{Y}$ is an internal factor of $XY$ (a contradiction with the hypothesis), or $Y$ is an internal factor of $YY$ (which contradicts the fact that $Y$ is primitive). Thus $Z$ must be empty.
It follows that $B=YU$, hence $B=Y^{h}$, since $B$ is a suffix of $XY^{j}$. If we have $h\geq 2$, then $j\geq 2$, and since $B$ is a prefix of $XY^{j}$, and $X$ is nonempty and shorter that $Y$, we see that $Y$ is an internal factor of $YY$, a contradiction again. Thus $h=1$. ∎
@@
and, as $V_{m-1}$ is a suffix of $V_{m}$, we get, by comparing suffixes of the same length, that $V_{m-1}=V_{m-3}V_{m-2}^{b_{m-1}}$. Since also $V_{m-1}=V_{m-2}V_{m-3}V_{m-2}^{b_{m-1}-1}$, we deduce that $V_{m-3}V_{m-2}=V_{m-2}V_{m-3}$, a contradiction. Thus $b_{m-1}-d_{m-1}=0$.
@@
Suppose that the representation (8.2) is greedy. If $d_{m}=b_{m}$, then $V_{m}$ is not a suffix of $V_{m}V_{m-1}$.
Proof.
@@
To conclude, note that $V_{m-2}$ is a prefix and a suffix of $V_{m}$, contradicting the fact that a Christoffel word has no border. ∎
@@
Observe that $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$ is a prefix of $V_{m}=V_{m-2}V_{m-2}^{b_{m-1}}V_{m-3}\cdots$ if and only if $V_{m-3}$ is a prefix of $V_{m-2}$, thus, by Corollary 8.9, if and only if $V_{m-2}$ is not the Christoffel word $V_{m-2}(\ldots,0,b_{m-2})$. But $V_{m-2}$ cannot be equal to the latter word, since $V_{m}$ is by assumption not a Christoffel word, and $d_{m}=b_{m},d_{m-1}=0$. Thus $V_{m-1}$ is a prefix of $V_{m}$, and $B=V_{m-1}$ by Lemmas 8.13 (applied to m1m-1) and 8.16.

### v5

@@ 8 Borders of conjugates of Christoffel words
8.1 Borders
In this subsection, we determine the longest border of every conjugate of a Christoffel word, thereby reproving a result of Lapointe Lapointe (2017) (but with a totally different method). Indeed, the length of the longest border and the smallest nontrivial period of a word are related: their sum is the length of the word.
@@ 8.1 Borders
The key point for the proof of Theorem 8.1 is the determination of all the occurrences of $V_{m-1}$ in $V_{m}$. Exactly $b_{m}$ of them can be read on the factorization (15), but there may be additional ones. By primitivity of $V_{m-1}$, the word $V_{m-1}V_{m-1}$ contains exactly two occurrences of $V_{m-1}$. Consequently, if an additional occurrence of $V_{m-1}$ appears, then it must be a factor of $V_{m-1}V_{m-2}V_{m-1}$. A more precise statement is given in Lemma 8.25.
@@ 8.2 Consequences
As noted by one of the referees, this result may also be obtained as follows: if $u$ is a border of the Christoffel word $w$, then $uu$ is a factor of $ww$; since $ww$ is a Sturmian word, it is Sturmian, and thus $uu$ too. Hence all conjugates of $u$ are Sturmian. Among them is the power $\ell^{k}$ of some Lyndon word $\ell$, which is therefore Sturmian; hence $\ell$ is a Christoffel word by a theorem of Berstel and de Luca (see Reutenauer (2019) Corollary 13.4.3).
@@
where $M_{m}=V_{m}(0,\ldots,0)$ is as before the word corresponding to the sequence $a_{1},\ldots,a_{m}$ and to the Ostrowski representation of 0. Note that here $N$ may be in $\mathbb{Z}$; but the word $H_{N}(a_{1},\ldots,a_{m})$ depends only on $N$ modulo $q_{m}$, where $q_{m}$ is the length of this word (recall that $q_{j}=K_{j}(a_{1},\ldots,a_{j})$).
@@
Note that we recover a step function (the number $t$ in the statement) as it appears in Lapointe’s article, see for example (Lapointe, 2017, Figure 3).
@@
Proof 8.5.
Denote the new sequence by $a^{\prime}_{1}=a_{2}+1,a^{\prime}_{2}=a_{3},\ldots,a^{\prime}_{m-1}=a_{m}$; then the associated sequence of $b_{i}$’s is $b^{\prime}_{1}=a^{\prime}_{1}-1=a_{2},b^{\prime}_{2}=a^{\prime}_{2},\ldots,b^{\prime}_{m-1}=a^{\prime}_{m-1}$. Recall that $V_{-1}=V^{\prime}_{-1}=b,V_{0}=V^{\prime}_{0}=a$. Observe that $V_{1}=b$, $V_{2}=b^{a_{2}-d_{2}}ab^{d_{2}}$, while $V^{\prime}_{1}=a^{a_{2}-d_{2}}ba^{d_{2}}$. Thus, $V_{1}=E(V^{\prime}_{0})$ and $V_{2}=E(V^{\prime}_{1})$. An immediate induction based on (11) proves the lemma.
@@
Proof 8.7.
Let $q_{i}$ be as usual and write $q^{\prime}_{i}$ for the corresponding numbers with respect to the sequence $a_{1},\ldots,a_{m-2},a_{m-1}+1$. By stability, we have $V_{i}=V^{\prime}_{i}$ for $i=1,\ldots,m-2$. Next, we see that $V_{m-1}=V_{m-2}^{b_{m-1}-d_{m-1}}V_{m-3}V_{m-2}^{d_{m-1}}$. Thus,
$V_{m}(d_{1},\ldots,d_{m-1},0)=V_{m-1}V_{m-2}=V_{m-2}^{b_{m-1}-d_{m-1}}V_{m-3}V_{m-2}^{d_{m-1}+1}\\ ={V^{\prime}}_{m-2}^{b_{m-1}-d_{m-1}}V^{\prime}_{m-3}{V^{\prime}}_{m-2}^{d_{m-1}+1}$
$V^{\prime}_{m-1}(d_{1},\ldots,d_{m-2},d_{m-1}+1).$
@@
${V^{\prime}}_{m-2}^{b_{m-1}-d_{m-1}+1}V^{\prime}_{m-3}{V^{\prime}}_{m-2}^{d_{m-1}}$
$V^{\prime}_{m-1}(d_{1},\ldots,d_{m-2},d_{m-1}),$
and the proof is complete.
Proof 8.8 (of Corollary 8.3).
@@
2. We study now the number $t$ in Part (a) of the statement. Since $N=d_{m}q_{m-1}+\sum_{1\leq i\leq m-1}d_{i}q_{i-1}$, we get $\lfloor\frac{N}{q_{m-1}}\rfloor=d_{m}$ by Lemma 10.3 (i).
@@
We show that the numerator in the latter fraction is always in the interval $q_{m-1},q_{m-1})$, so that the integer part of this fraction is either -1 or 0, and we give the condition when it is 0. We have indeed $q_{m-1}\leq q_{m-2}-2-j$, since $j\leq q_{m-1}-1$ by Lemma 10.3 (i), so that $j+2\leq q_{m-1}+1\leq q_{m-1}+q_{m-2}$. Moreover, $q_{m-2}-2-j<q_{m-1}$, since the sequence $q_{i}$ is increasing. Also, if $d_{m-1}>0$, then $j\geq q_{m-2}$, hence $q_{m-2}-2-j<0$; and if $d_{m-1}=0$, then $j=\sum_{1\leq i\leq m-2}d_{i}q_{i-1}$, and $q_{m-2}-2-j<0$ if and only if $j\geq q_{m-2}-1$, which, by Lemmas 10.1 (i) and 10.3 (i), is equivalent to the fact that the sequence $d_{1},\ldots,d_{m-2},0$ is alternating.
@@
Suppose now that $q_{m-1}\leq N<q_{m}$. Then $d_{m}>0$ by Lemma 10.3 (i). We are therefore in case (i), (ii), (iii) or (iv) of Theorem 8.1. Note that $V_{m-1}=H_{J}(a_{1},\ldots,a_{m-1})$, where $J=\sum_{1\leq i\leq m-1}d_{i}q_{i-1}$, so that $V_{m-1}=H_{N}(a_{1},\ldots,a_{m-1})$, because $N$ is congruent to $J$ modulo $q_{m-1}$, the length of $V_{m-1}$. Thus we have to show that $B_{N}=V_{m-1}^{t}$.
@@
5. We assume that $a_{m}=1$. Define $a^{\prime}_{i}=a_{i}$ if $i=1,\ldots,m-2$ and $a^{\prime}_{m-1}=a_{m-1}+1$. We denote by $q^{\prime}_{i}$ and $M^{\prime}_{i}$ the corresponding words and numbers. We have $q^{\prime}_{i}=q_{i}$ for $i=1,\ldots,m-2$ and by (2), $q^{\prime}_{m-1}=q_{m-1}+q_{m-2}=q_{m}$.
@@
Suppose that $0\leq N<q_{m-1}-1$. Then $q^{\prime}_{m-2}=q_{m-2}\leq N^{\prime}<q_{m-1}+q_{m-2}-1=q_{m}-1=q^{\prime}_{m-1}-1$. It follows from Case (a) that $B_{N}=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})^{t}$, with $t=\min\{\lfloor\frac{N^{\prime}}{q^{\prime}_{m-2}}\rfloor,1+\lfloor\frac{q^{\prime}_{m-1}-2-N^{\prime}}{q^{\prime}_{m-2}}\rfloor\}$. Note that we have $H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2})=C^{N^{\prime}}(H_{0}(a^{\prime}_{1},\ldots,a^{\prime}_{m-2}))=C^{N+q_{m-2}}(H_{0}(a_{1},\ldots,a_{m-2}))$ $C^{N}(H_{0}(a_{1},\ldots,a_{m-2}))$ (since the word is of length $q_{m-2}$) $H_{N}(a_{1},\ldots,a_{m-2})$; moreover, $t=\min\{\lfloor\frac{N+q_{m-2}}{q_{m-2}}\rfloor,1+\lfloor\frac{q_{m-1}+q_{m-2}-2-N-q_{m-2}}{q_{m-2}}\rfloor\}$, which settles this case.
Suppose now that $q_{m-1}\leq N<q_{m}-1$. Then we have $H_{N}(a_{1},\ldots,a_{m})=H_{N^{\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$ $H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1})$, where $N^{\prime\prime}=N^{\prime}-q_{m}=N+q_{m-2}-q_{m}=N-q_{m-1}$, since the words have length $q_{m}$. Now $0\leq N^{\prime\prime}<q_{m-2}-1$. Hence by the first part, $B_{N}=H_{N^{\prime\prime}}(a^{\prime}_{1},\ldots,a^{\prime}_{m-1}-1)=C^{N^{\prime\prime}}(H_{0}(a_{1},\ldots,a_{m-1}))=C^{N}(H_{0}(a_{1},\ldots,a_{m-1}))=H_{N}(a_{0},\ldots,a_{m-1})$, since the word has length $q_{m-1}$.
8.3 Proof of Theorem 8.1
@@
Since $q_{j}$ is the length of $V_{j}$, this proves the lemma.
@@
Proof 8.12.
The expression of $X_{k}W_{k}$ is an immediate consequence of Lemma 8.9. Recall Pirillo’s theorem: if the words $aub,bua$ are conjugate, then $u$ is a central word (Pirillo (1999), (Reutenauer, 2019, Theorem 15.2.5)). By Lemma 8.9, the words $aX_{k}W_{k}b$ and $bX_{k}W_{k}a$ are conjugate. This proves the lemma.
@@
Proof 8.14.
If $V_{k+1}$ is a prefix of $V_{k}V_{k+1}$, then the common prefix $W_{k}$ of $V_{k}V_{k+1}$ and $V_{k+1}V_{k}$ is of length $w_{k}\geq q_{k+1}$; since $w_{k}+x_{k}=q_{k+1}+q_{k}-2$, we obtain $x_{k}\leq q_{k}-2$. By Lemma 8.9 this implies that $d_{k+1}=0$. Similarly, if $V_{k+1}$ is a suffix of $V_{k+1}V_{k}$, then $w_{k}\leq q_{k}-2$ and $b_{k+1}-d_{k+1}=0$.
@@
Suppose that the representation (16) is greedy. The word $V_{m-1}(d_{1},\ldots,d_{m-1})$ is a prefix of $V_{m}(d_{1},\ldots,d_{m})$ if and only if $V_{m}(d_{1},\ldots,d_{m})$ is not the Christoffel word $V_{m}(\ldots,0,b_{m-2},0,b_{m})$.
Proof 8.16.
@@
The lemma then follows from Lemma 3.2 and Corollary 7.7.
@@
Suppose that the representation (16) is greedy. The word $V_{m-1}$ is not a prefix of the word $V_{m-2}V_{m-1}$ if and only if $d_{m-1}=0$ and the sequence $d_{1},\ldots,d_{m-1}$ is alternating.
Proof 8.18.
We apply Corollary 8.15 to the sequence $a_{1},\ldots,a_{m-1},1$ and the words $V_{m-1}(d_{1},\ldots,d_{m-1})$ and $V^{\prime}_{m}=V_{m}(d_{1},\ldots,d_{m-1},1)$, so that $V^{\prime}_{m}=V_{m-2}V_{m-1}$: thus the word $V_{m-1}$ is a prefix of $V_{m-2}V_{m-1}$ if and only if $V^{\prime}_{m}$ is not a Christoffel word; but, by Corollary 7.7, $V^{\prime}_{m}$ is a Christoffel word if and only if the sequence $d_{1},\ldots,d_{m-1},1$ is alternating; this means that $d_{m-1}=0$ and the sequence $d_{1},\ldots,d_{m-1}$ is alternating.
@@
Proof 8.20.
The word $Y^{i-1}X$ is a border of $Y^{i}X$. Suppose that $B$ is longer. It begins by $Y^{i-1}X$, hence by $Y$ since $i\geq 2$: $B=YB^{\prime}$. Moreover, $Y^{i}X=UB$, where the length of $U$ satisfies $0<|U|<|Y|$. Thus $Y^{i}X=UYB^{\prime}$ and we see that $Y$ is an internal factor of $YY$, a contradiction.
@@
(ii) Follows from (i) and Lemma 8.19.
@@
In the following lemmas, we consider the legal representation (16) and put $V_{m}=V_{m}(d_{1},\ldots,d_{m})$, $V_{m-1}=V_{m-1}(d_{1},\ldots,d_{m-1})$, and $V_{m-2}=V_{m-2}(d_{1},\ldots,d_{m-2})$.
@@
The proof for $V_{m-1}V_{m}$ is similar and we omit it.
@@
1. Assume first that $1\leq d_{m}\leq b_{m}-1$. Thus by (15), the word $V_{m-1}$ is a prefix and a suffix of $V_{m}$. We show that either $X$ is an integer power of $V_{m-1}$, or $Y$ is an integer power of $V_{m-1}$. Indeed, if $|X|<|V_{m-1}|$, then $X$ is a nontrivial proper suffix of $V_{m-1}$, $V_{m-1}=UX$, where $U$ is nonempty, and $V_{m}$ begins with $XV_{m-1}$; but $V_{m}$ also begins with $V_{m-1}$, $V_{m}=V_{m-1}W$, hence $UV_{m-1}W=UV_{m}=UXV_{m-1}Y=V_{m-1}V_{m-1}Y$; since $U$ is nonempty and shorter than $V_{m-1}$, we see that $V_{m-1}$ is a proper factor of $V_{m-1}^{2}$, and we have a contradiction with the primitivity of $V_{m-1}$.
@@
Suppose now that there is an occurrence of $V_{m}$ ending at $V_{m-1}$; this is equivalent to the fact that $V_{m}$ is a suffix of $V_{m}V_{m-1}$. Then, similarly, we must have $d_{m}=b_{m}$.
@@
Suppose that $B$ is longer; then $B$ extends $Y^{i}$ to the left, hence $B$ ends by $Y$, since $i\geq 1$; moreover, we may extend the prefix $Y^{i}$ of $W$ to the longer prefix $B$, and since $B$ ends by $Y$, by 1. this $Y$ is the factor $Y$ of $W$ starting at $X$, and we are in case L; thus, since $B$ ends by $Y$, by 1. and by the hypothesis on the locations of the factors $Y$ in $W$, $B=Y^{i+1}$. But $B$ is also a right factor of $W$, hence we must have $j>i$, otherwise there is a factor $Y$ ending at $X$, which is excluded by 2.
@@
Then, since $B$ is a suffix of $XY^{j}$, we have $XY^{j}=W{\widetilde{Y}}UZ$; then we see that that either $\widetilde{Y}$ is an internal factor of $XY$ (a contradiction with the hypothesis), or $\widetilde{Y}$ is an internal factor of $YY$ (which contradicts the fact that $Y$ is primitive). Thus $Z$ must be empty.
It follows that $B=YU$, hence $B=Y^{h}$, since $B$ is a suffix of $XY^{j}$. If we have $h\geq 2$, then $j\geq 2$, and since $B$ is a prefix of $XY^{j}$, and $X$ is nonempty and shorter that $Y$, we see that $Y$ is an internal factor of $YY$, a contradiction again. Thus $h=1$.
@@
and, as $V_{m-1}$ is a suffix of $V_{m}$, we get, by comparing suffixes of the same length, the equality $V_{m-1}=V_{m-3}V_{m-2}^{b_{m-1}}$. Since also $V_{m-1}=V_{m-2}V_{m-3}V_{m-2}^{b_{m-1}-1}$, we deduce that $V_{m-3}V_{m-2}=V_{m-2}V_{m-3}$, a contradiction. Thus $b_{m-1}-d_{m-1}=0$.
@@
Suppose that the representation (16) is greedy. If $d_{m}=b_{m}$, then $V_{m}$ is not a suffix of $V_{m}V_{m-1}$.
Proof 8.34.
@@
To conclude, note that $V_{m-2}$ is a prefix and a suffix of $V_{m}$, contradicting the fact that a Christoffel word has no border.
@@
Observe that $V_{m-1}=V_{m-2}^{b_{m-1}}V_{m-3}$ is a prefix of $V_{m}=V_{m-2}V_{m-2}^{b_{m-1}}V_{m-3}\cdots$ if and only if $V_{m-3}$ is a prefix of $V_{m-2}$, thus, by Corollary 8.15, if and only if $V_{m-2}$ is not the Christoffel word $V_{m-2}(\ldots,0,b_{m-2})$. But $V_{m-2}$ cannot be equal to the latter word, since $V_{m}$ is by assumption not a Christoffel word, and $d_{m}=b_{m},d_{m-1}=0$. Thus $V_{m-1}$ is a prefix of $V_{m}$, and $B=V_{m-1}$ by Lemmas 8.23 (applied to m1m-1) and 8.29.